OpenStudy (anonymous):

Multiply x¾ X ⅛ A. x^1/8 B. x^3/10 C. x^11/12 D. x^2/5 Anyone? i'm not sure if i'm supposed to multiply the fractions or cross multiply..?

OpenStudy (anonymous):

anyonee?

OpenStudy (anonymous):

Powers to powers, multiply. when multiplying a set of powers you add.

OpenStudy (anonymous):

That just completely confused me. So, i'm supposed to add them? i don't get any of the answers given though.

OpenStudy (anonymous):

you add them, think of taking 5^(1/2) * 5^(1/2) aka the square root times the square root, you know that two square roots multiplied gives you back just 5. So then think do you add 1/2+1/2 to get 5^1 or multiply. because 5^(1/4) doesn't equal 5. then you know it's based on adding

OpenStudy (anonymous):

the answer should be x^(7/8) idk why it's not in your choices

OpenStudy (anonymous):

Well those are the only answers i am given, so is it possible to do this another way?

OpenStudy (anonymous):

unfortunately no. it's a math rule. you can try it with other numbers and prove it's adding. 14^(1/4)* 14^(4/11) = 5.050168167 also written 14^(27/44)

OpenStudy (anonymous):

OH. wait a second, instead of the ^1/8 it's ^1/6

OpenStudy (anonymous):

then it's x^11/12

OpenStudy (anonymous):

So the answer is x^1/8 correct?

OpenStudy (anonymous):

oh, okay. you sure?

OpenStudy (anonymous):

yes it's x^(11/12)

OpenStudy (anonymous):

Okay, thank youu. can you help me with one more? i'll post it in another question

OpenStudy (anonymous):

Look. These are all the Exponent Laws you need to know: \[A^x \times A^y=A^{x+y}\]\[\frac{ A^x }{ A^y }=A^{x-y}|A^y \ne 0\]\[(A^x)^y=A^{xy}\]\[A^xB^x=(AB)^x\]\[\frac{ A^x }{ B^x }=\left( \frac{ A }{ B } \right)^x\]\[A^0=1|A \ne 0\]\[A^{-x}=\frac{ 1 }{ A^x }|A^x \ne 0\]You're question is x^(3/4) * x^(1/8). To solve this, we use the Product Exponent Law, which is the first one in this list. Using it we get:\[A^x \times A^y=A^{x+y} \implies x^{3/4} \times x^{1/8} =x^{(3/4)+(1/8)}=x^{7/8}\]Get it? @brazukinhaa

OpenStudy (anonymous):

es, and i got 4/10 but that's not an option.. And instead of 1/8 it's 1/6